OpenStudy (anonymous):

Find the first TWO non-negative asymptotes and the first negative asymptote of the graph of y in radians y=tan(3x) smallest non-negative asymptote: x= second non-negative asymptote: x= first negative asymptote: x=

OpenStudy (anonymous):

smallest non-negative asymptote: x= pi/6 second non-negative asymptote: x= pi/6 + pi/3 = pi/2 first negative asymptote: x= -pi/6 inflection point (on the midline at x=0 period = pi/3

OpenStudy (anonymous):

omg thanks can you help me with more of the same type of questions pls

OpenStudy (anonymous):

I'll try

OpenStudy (anonymous):

ok let me type em up

OpenStudy (anonymous):

heres another one Find the first TWO non-negative asymptotes and the first negative asymptote of the graph of y in radians y=cot(x−π/3) smallest non-negative asymptote: x= second non-negative asymptote: x= first negative asymptote: x=

OpenStudy (anonymous):

period = pi smallest non-negative asymptote: x= pi/3 (this is the shift to the right) second non-negative asymptote: x= pi/3 + pi = 4pi/3 first negative asymptote: x= pi/3 - pi = 2pi/3

OpenStudy (anonymous):

wow man thanks for the help. i have 4 more of these types of question...do you have the time man? i dont want to bother you

OpenStudy (anonymous):

let's go

OpenStudy (anonymous):

Find the first TWO non-negative asymptotes and the first negative asymptote of the graph of y in radians y=sec(3x). smallest non-negative asymptote: x= second non-negative asymptote: x= first negative asymptote: x=

OpenStudy (anonymous):

period =2pi/3 the asymptotes will occur where the cos[3x] is equal to 0: smallest non-negative asymptote: x= pi/6 second non-negative asymptote: x= pi/2 first negative asymptote: x= -pi/6

OpenStudy (anonymous):

Find the first TWO non-negative asymptotes and the first negative asymptote of the graph of y in radians y=csc(x/5). smallest non-negative asymptote: x= second non-negative asymptote: x= first negative asymptote: x=

OpenStudy (anonymous):

period = 10pi there is no horizontal shift so smallest non-negative asymptote: x= 0 second non-negative asymptote: x= 5pi (half way from 0 to 10pi) first negative asymptote: x= -5pi (half way from 0 to -10pi)

OpenStudy (anonymous):

Find the first TWO non-negative asymptotes and the first negative asymptote of the graph of y in radians y=2sec(x−π/3)+1. smallest non-negative asymptote: x= second non-negative asymptote: x= first negative asymptote: x=

OpenStudy (anonymous):

y=2sec(x−π/3)+1. The 2 and the +1 do not have anything to do with the asymptotes period = 2pi smallest non-negative asymptote: x= 5pi/6 second non-negative asymptote: x= 11pi/6 first negative asymptote: x= -pi/6

OpenStudy (anonymous):

Find the first TWO non-negative asymptotes and the first negative asymptote of the graph of y in radians y=−8csc(5x+π/8)−2. smallest non-negative asymptote: x= second non-negative asymptote: x= first negative asymptote: x=

OpenStudy (anonymous):

Sorry, I've gtg for a short while so this is the last one I can help you with or now. y=−8csc(5x+π/8)−2. period = 2pi/5 half period =pi/5 NOTE ORDER OF ANSWERS CHANGED... first negative asymptote: x= -pi/8 (horizontal shift left) smallest non-negative asymptote: x= -pi/8 + pi/5 = 3pi/40 second non-negative asymptote: x= 3pi/40 + pi/5 = 11pi/40

OpenStudy (anonymous):

yea none none worked

OpenStudy (anonymous):

its -9csc(5x+pi/9)-2

OpenStudy (anonymous):

help please?